# The Hidden Life of Tokens: Reducing Hallucination of Large Vision-Language Models via Visual Information Steering

### v1

@@ Abstract
Large Vision-Language Models (LVLMs) can reason effectively over both textual and visual inputs, but they tend to hallucinate syntactically coherent yet visually ungrounded contents. In this paper, we investigate the internal dynamics of hallucination by examining the tokens logits ranking throughout the generation process, revealing three key patterns in how LVLMs process information: (1) gradual visual information loss – visually grounded tokens gradually become less favored throughout generation, and (2) early excitation – semantically meaningful tokens achieve peak activation in the layers earlier than the final layer. (3) hidden genuine information – visually grounded tokens though not being eventually decoded still retain relatively high rankings at inference. Based on these insights, we propose VISTA (Visual Information Steering with Token-logit Augmentation), a training-free inference-time intervention framework that reduces hallucination while promoting genuine information. VISTA works by combining two complementary approaches: reinforcing visual information in activation space and leveraging early layer activations to promote semantically meaningful decoding. Compared to existing methods, VISTA requires no external supervision and is applicable to various decoding strategies. Extensive experiments show that VISTA on average reduces hallucination by about 40% on evaluated open-ended generation task, and it consistently outperforms existing methods on four benchmarks across four architectures under three decoding strategies. Code is available at: https://github.com/LzVv123456/VISTA
@@ 1 Introduction
Though LVLM hallucination is considered multifaceted (Liu et al., 2024d), a critical cause stems from the overwhelming influence of language priors over visual contexts, and has been studied from the perspective of attention patterns (Huang et al., 2024; Liu et al., 2024f) and distribution divergence within logits space (Leng et al., 2024; Favero et al., 2024). Despite these insights, it remains unclear how hallucination emerges and propagates during the generation process.
Inspecting Token Dynamics in LVLMs. In this work, we take a novel perspective by examining LVLM’s generation dynamics through the lens of token logits ranking. Given an image its corresponding description (produced by an LVLM), we identify three categories of tokens (elaborated in Sec. 3.4):
@@ 2 Methodology
In this section, we first establish the notions and knowledge foundation in Sec. 2.1, followed by the elaboration of token ranking analysis in Sec. 2.2. We then present the proposed VSV and SLA methods in Sec. 2.3 and 2.4, respectively.
@@ 2.3 Visual Steering Vector (VSV)
Being aware of the challenge from gradual visual information loss, it is of critical importance to retain visual cues throughout the generation. A promising method involves increasing attention weights distributed on visual tokens (Liu et al., 2024f). Nevertheless, this operation simultaneously introduces undesired parametric priors cumulated in residual streams of visual tokens. Drawing inspiration from steering vectors in LLMs (Turner et al., 2023; Zou et al., 2023; Liu et al., 2024e; Li et al., 2024), we propose Visual Steering Vector (VSV) to steer the generation of LVLM towards the direction with visual groundings without amplifying inherent language biases.
@@ 3.1 Experimental Setup
Baselines. Besides three vanilla decoding strategies, we set aside VISTA with several SoTA hallucination mitigation methods that can operate without external supervision. DoLa (Chuang et al., 2024) signifies a internal (across layers) contrastive strategy within logits space, while VCD (Leng et al., 2024) presents a parallel contrastive method across time steps. OPERA (Huang et al., 2024) is a powerful technique tailored for the beam search. PAI (Liu et al., 2024f) is another inference-time intervention method that is most comparable to ours. We reproduce all baseline results using identical evaluation data and settings (e.g., prompt, temperature). Methods without official supports for certain architectures and decoding strategies are omitted to prevent implementation bias.
@@ 3.2 Results on Object Hallucination Benchmarks
CHAIR Evaluation. Caption Hallucination Assessment with Image Relevance (CHAIR) (Rohrbach et al., 2018) provides a systematic framework for evaluating object hallucination in image captioning tasks. CHAIR assesses caption accuracy by comparing mentioned objects against ground-truth labels, with hallucinations defined as objects present in captions but absent from ground truth. The metric operates at two levels: instance-level ($\text{CHAIR}_{\text{I}}$) and sentence-level ($\text{CHAIR}_{\text{S}}$): $\text{CHAIR}_{I}=\frac{\left|\{\text{hallucinated object}\}\right|}{\left|\{\text{object}\}\right|}$ and $\text{CHAIR}_{S}=\frac{\left|\{\text{caption w/ hallucinated objects}\}\right|}{\left|\{\text{caption}\}\right|}.$ Following established protocol (Huang et al., 2024; Liu et al., 2024f), we evaluate on 500 randomly sampled images from MSCOCO 2014 validation set, using the prompt “Please help me describe the image in detail” with maximum generation length of 512 tokens.
Results in Table 1 show that VISTA significantly reduces hallucination in open-ended generation task, outperforming existing inference-time intervention method and other contrastive decoding methods by a substantial margin. VISTA brings around 40% relative improvement upon corresponding vanilla decoding methods. Notably, while PAI (Liu et al., 2024f) shows less efficacy for sampling-based decoding, VISTA excels across all decoding strategies. We attribute this robust performance to the contrastive design and the choice of activation space steering which is not hinged with any specific decoding strategy.
@@ 4 Related Work
Hallucination Mitigation in LVLMs. Hallucination – the generation of content that is irrelevant, factually incorrect, or inconsistent with visual inputs (Bai et al., 2024) – represents a fundamental challenge in LVLM development. Research has identified three primary sources: limitations in visual encoder capabilities (Tong et al., 2024; Liu et al., 2024b; Shi et al., 2024), excessive reliance on learned parametric knowledge (Li et al., 2023b; Zhou et al., 2023; Leng et al., 2024; Huang et al., 2024), and noisy training data (Liu et al., 2023; Yu et al., 2024). Mitigation approaches span training-based solutions with refined datasets (Yue et al., 2024; Jiang et al., 2024), post-processing techniques including revision (Yin et al., 2023; Zhou et al., 2023) and verification (Chen et al., 2024; Sun et al., 2023), and inference-time interventions like Visual Contrastive Decoding (Leng et al., 2024) and enhanced attention methods (Liu et al., 2024f). Recent studies revealing “text inertia” (Liu et al., 2024f), where models generate similar hallucinations without visual input, highlight concerning reliance on learned text patterns. While these findings advance our understanding, how hallucination propagates through model architectures remains elusive, and existing solutions often require external supervision and are hinged with specific decoding strategies.
@@ A.3 Advantages of Token Ranking Analysis
Why Not Attention? While attention matrices have been extensively studied in LVLM hallucination research to understand information flow patterns and inform mitigation strategies (Huang et al., 2024; Liu et al., 2024f), our token ranking methodology offers several distinct advantages. Previous work, such as PAI (Liu et al., 2024f), has attributed hallucination phenomena like “text inertia” to insufficient attention allocation to visual tokens. However, this interpretation is potentially confounded by the presence of “anchor tokens” (Wang et al., 2023; Huang et al., 2024) that aggregate and redistribute information across the network. The existence of these information hubs means that reduced attention weights on visual tokens, particularly in later layers where visual information can be accessed indirectly through anchor positions, may not necessarily indicate information loss.

### v2

@@ Abstract
Large Vision-Language Models (LVLMs) can reason effectively over both textual and visual inputs, but they tend to hallucinate syntactically coherent yet visually ungrounded contents. In this paper, we investigate the internal dynamics of hallucination by examining the tokens logits ranking throughout the generation process, revealing three key patterns in how LVLMs process information: (1) gradual visual information loss – visually grounded tokens gradually become less favored throughout generation, and (2) early excitation – semantically meaningful tokens achieve peak activation in the layers earlier than the final layer. (3) hidden genuine information – visually grounded tokens though not being eventually decoded still retain relatively high rankings at inference. Based on these insights, we propose VISTA (Visual Information Steering with Token-logit Augmentation), a training-free inference-time intervention framework that reduces hallucination while promoting genuine information. VISTA works by combining two complementary approaches: reinforcing visual information in activation space and leveraging early layer activations to promote semantically meaningful decoding. Compared to existing methods, VISTA requires no external supervision and is applicable to various decoding strategies. Extensive experiments show that VISTA on average reduces hallucination by about 40% on evaluated open-ended generation task, and it consistently outperforms existing methods on four benchmarks across four architectures under three decoding strategies. Code is available at https://github.com/LzVv123456/VISTA.
@@ 1 Introduction
Though LVLM hallucination is considered multifaceted (Liu et al., 2024d), a critical cause stems from the overwhelming influence of language priors over visual contexts, and has been studied from the perspective of attention patterns (Huang et al., 2024; Liu et al., 2024g) and distribution divergence within logits space (Leng et al., 2024; Favero et al., 2024). Despite these insights, it remains unclear how hallucination emerges and propagates internally during the course of generation.
Inspecting Token Dynamics in LVLMs. In this work, we take a novel perspective by examining LVLM’s generation dynamics through the lens of token logits ranking. Given an image and its corresponding description (produced by an LVLM), we identify three categories of tokens (elaborated in Sec. 3.4):
@@ 2 Methodology
In this section, we first establish the notions and knowledge foundation in Sec. 2.1, followed by the elaboration of token ranking analysis in Sec. 2.2. We then present the proposed VSV and SLA methods in Sec. 2.3 and Sec. 2.4, respectively.
@@ 2.3 Visual Steering Vector (VSV)
Being aware of the challenge from gradual visual information loss, it is of critical importance to retain visual cues throughout the generation. A promising method involves increasing attention weights distributed on visual tokens (Liu et al., 2024g). Nevertheless, this operation simultaneously introduces undesired parametric priors cumulated in residual streams of visual tokens. Drawing inspiration from steering vectors in LLMs (Turner et al., 2023; Zou et al., 2023; Liu et al., 2024e; Li et al., 2024), we propose Visual Steering Vector (VSV) to steer the generation of LVLM towards the direction with visual groundings without amplifying inherent language biases.
@@ 3.1 Experimental Setup
Baselines. Besides three vanilla decoding strategies, we compare VISTA with several SoTA hallucination mitigation methods that can operate without external supervision. DoLa (Chuang et al., 2024) signifies a internal (across layers) contrastive strategy within logits space, while VCD (Leng et al., 2024) presents a parallel contrastive method across time steps. OPERA (Huang et al., 2024) is a powerful technique tailored for the beam search. PAI (Liu et al., 2024g) is another inference-time intervention method that is most comparable to ours. We reproduce all baseline results using identical evaluation data and settings (e.g., prompt, temperature). Methods without official supports for certain architectures and decoding strategies are omitted to prevent implementation bias.
@@ 3.2 Results on Object Hallucination Benchmarks
CHAIR Evaluation. Caption Hallucination Assessment with Image Relevance (CHAIR) (Rohrbach et al., 2018) provides a systematic framework for evaluating object hallucination in image captioning tasks. CHAIR assesses caption accuracy by comparing mentioned objects against ground-truth labels, with hallucinations defined as objects present in captions but absent from ground truth. The metric operates at two levels: instance-level ($\text{CHAIR}_{\text{I}}$) and sentence-level ($\text{CHAIR}_{\text{S}}$): $\text{CHAIR}_{I}=\frac{\left|\{\text{hallucinated object}\}\right|}{\left|\{\text{object}\}\right|}$ and $\text{CHAIR}_{S}=\frac{\left|\{\text{caption w/ hallucinated objects}\}\right|}{\left|\{\text{caption}\}\right|}.$ Following established protocol (Huang et al., 2024; Liu et al., 2024g), we evaluate on 500 randomly sampled images from MSCOCO 2014 validation set, using the prompt “Please help me describe the image in detail” with maximum generation length of 512 tokens.
Results in Table 1 show that VISTA significantly reduces hallucination in open-ended generation task, outperforming existing inference-time intervention method and other contrastive decoding methods by a substantial margin. VISTA brings around 40% relative improvement upon corresponding vanilla decoding methods. Notably, while PAI (Liu et al., 2024g) shows less efficacy for sampling-based decoding, VISTA excels across all decoding strategies. We attribute this robust performance to the contrastive design and the choice of activation space steering which is not hinged with any specific decoding strategy.
@@ 4 Related Work
Hallucination Mitigation in LVLMs. Hallucination – the generation of content that is irrelevant, factually incorrect, or inconsistent with visual inputs (Bai et al., 2024) – represents a fundamental challenge in LVLM development. Research has identified three primary sources: limitations in visual encoder capabilities (Tong et al., 2024; Liu et al., 2024b; Shi et al., 2024), excessive reliance on learned parametric knowledge (Li et al., 2023b; Zhou et al., 2023; Leng et al., 2024; Huang et al., 2024), and noisy training data (Liu et al., 2023; Yu et al., 2024). Mitigation approaches span training-based solutions with refined datasets (Yue et al., 2024; Jiang et al., 2024), post-processing techniques including revision (Yin et al., 2023; Zhou et al., 2023) and verification (Chen et al., 2024; Sun et al., 2023), and inference-time interventions like Visual Contrastive Decoding (Leng et al., 2024) and enhanced attention methods (Liu et al., 2024g). Recent studies revealing “text inertia” (Liu et al., 2024g), where models generate similar hallucinations without visual input, highlight concerning reliance on learned text patterns. Similar to VISTA, VTI (Liu et al., 2024f) also applies representation engineering technique to LVLM hallucination domain, yet with a focus on the instability issue of visual representations. While prior findings advance our understanding, how hallucination propagates through model architectures remains elusive, and existing solutions often require external supervision or are hinged with specific decoding strategies.
@@ A.3 Advantages of Token Ranking Analysis
Why Not Attention? While attention matrices have been extensively studied in LVLM hallucination research to understand information flow patterns and inform mitigation strategies (Huang et al., 2024; Liu et al., 2024g), our token ranking methodology offers several distinct advantages. Previous work, such as PAI (Liu et al., 2024g), has attributed hallucination phenomena like “text inertia” to insufficient attention allocation to visual tokens. However, this interpretation is potentially confounded by the presence of “anchor tokens” (Wang et al., 2023; Huang et al., 2024) that aggregate and redistribute information across the network. The existence of these information hubs means that reduced attention weights on visual tokens, particularly in later layers where visual information can be accessed indirectly through anchor positions, may not necessarily indicate information loss.